# Bimetric MOND as a framework for variable-$G$ theories – local systems and cosmology

### v1

@@ IV Einstein-Palatini formulation
The Einstein-Palatini formulation of GR starts from the Einstein-Hilbert action $\propto\int g^{1/2}g^{\scriptscriptstyle\mu\nu}R_{\mu\nu}$, where the curvature is derived from a connection ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$, symmetric in $\mu\nu$. In the standard, “metric”, formulation of GR, the connection is taken as the Levi-Civita connection, $\Gamma^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$, derived from the metric – thus making the metric the only gravitational DoF. In the Einstein-Palatini formulation of GR, ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$, as it appears in the expression for the curvature scalar, is taken as an independent gravitational DoF. [^7] A connection appears also in the matter action, via covariant derivatives; there, it is standardly taken as the Levi-Civita, metric connection. Then, the field equations gotten by varying the action over ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$ tell us that for solutions of the theory we do have ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}=\Gamma^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$. Then, variation over the metric gives the standard Einstein equations for the metric, resulting in a theory equivalent to GR.
@@ IV.1 Field equation in the Einstein-Palatini formulation
Expression (31) can be viewed as the (uncoupled) Einstein-Hilbert action for the metrics $\mathfrak{g}_{\scriptscriptstyle\mu\nu}$ and $\hat{\mathfrak{g}}_{\scriptscriptstyle\mu\nu}$. However, this does not point to a simple formulation of the theory in terms of these metrics. Expressing the matter actions, and the BIMOND interaction Lagrangian density, $\tilde{\mathcal{M}}$, in terms of these metrics is highly involved, as matter couples minimally to $g_{\scriptscriptstyle\mu\nu}$ and $\hat{g}_{\scriptscriptstyle\mu\nu}$, not to $\mathfrak{g}_{\scriptscriptstyle\mu\nu}$ and $\hat{\mathfrak{g}}_{\scriptscriptstyle\mu\nu}$. This remains so, even when dealing with matter actions that are conformally invariant, such as the pure electromagnetic case, where the Maxwell action takes its usual form with $\mathfrak{g}_{\scriptscriptstyle\mu\nu}$ and $\hat{\mathfrak{g}}_{\scriptscriptstyle\mu\nu}$. As seen, e.g., from Eq. (34), even in this case, the presence of the BIMOND interaction is still an obstacle to a formulation in term of $\mathfrak{g}_{\scriptscriptstyle\mu\nu}$ and $\hat{\mathfrak{g}}_{\scriptscriptstyle\mu\nu}$.
@@ V Observational constraints on $\mathcal{M}_{\scriptscriptstyle G}$ and some examples
Before offering specific examples, I discuss what we require from such theories to achieve in the subcosmological realm, and the resulting constraints on the behaviour of $\mathcal{M}_{\scriptscriptstyle G}$.
V.0.1 Nonrelativistic limit
@@
These facts are used in Sec. V.0.4 to demonstrate that, with appropriate forms of $\mathcal{M}_{\scriptscriptstyle G}$, the various observational constraints can be easily satisfied.
V.0.2 High-acceleration limit
@@
This requirement applies to NR systems - such as the inner solar system, and noncompact stars - as well as to relativistic ones – such as black holes, pulsars, and merging compact objects. To achieve it, $\mathcal{M}_{\scriptscriptstyle G}$ has to depend also on variables that do not vanish in the NR limit – variables that are of the first type, but not of the second. We can then make $\mathcal{M}_{\scriptscriptstyle G}$ vanish rapidly as these variables tend to infinity, independently of the values of the variables of the second type on which $\mathcal{M}_{\scriptscriptstyle G}$ depends.
V.0.3 Low-acceleration limit
@@
V.0.4 Example
To summarize, I want to demonstrate that $\mathcal{M}_{\scriptscriptstyle G}$ can be chosen so that the following requirements are satisfied: (1) $\mathcal{M}_{\scriptscriptstyle G}$ vanishes rapidly for all high-acceleration systems -relativistic and NR, to comply with the many constraints from subcosmological systems on anomalous $G$ values. (2) $\mathcal{M}_{\scriptscriptstyle G}$ also vanishes rapidly in the NR limit even for systems that are not high-acceleration, so that pure-BIMOND phenomenology is predicted for galactic systems. (3) $\mathcal{M}_{\scriptscriptstyle G}$ can take finite values in later-time cosmology, while it vanishes for early-time cosmology. In this section I discuss the fulfilment of requirements (1) and (2) for subcosmological systems. Cosmology will be discussed in Sec. VI.
@@ V.0.4 Example
As we saw in Sec. V.0.1, for relativistic, subcosmological systems, of mass $M$, and characteristic size $\ell$, both the $X$ and $Y$ variables are of order $(MG/\ell^{2}a_{0})^{2}\ggg 1$ (because they are perforce of high accelerations). [^10] For highly relativistic systems, such as the vicinity or a black-hole horizon, $X,Y\sim\ell_{\scriptscriptstyle U}/\ell$, where $\ell_{\scriptscriptstyle U}$ is the Hubble distance. For NR, subcosmological systems, $Y=O(g^{2}/a_{0}^{2})$, where $g=MG/\ell^{2}$ for high-acceleration systems, and $g=(MGa_{0}/\ell^{2})^{1/2}$ for low-acceleration ones. And, $X=O(Y\epsilon^{2})$. In the low-acceleration case ($g\not\gg a_{0}$), we have $X\lesssim\epsilon^{2}\lll 1$. To ensure that requirement (2) is satisfied we need $\mathcal{M}_{\scriptscriptstyle G}$ to vanish rapidly enough in the low-acceleration case, as $X\lesssim\epsilon^{2}\rightarrow 0$.
To see how fast, consider the NR limit of the Einstein-Hilbert Lagrangian. We saw in Eqs. (29-30) that the post-Newtonically dominant term is $\vec{\nabla}\mathcal{M}_{\scriptscriptstyle G}\cdot\vec{\nabla}\epsilon$, which we want to be negligible compared with the Poissonian term $(\vec{\nabla}\epsilon)^{2}$. Since $X\lesssim\epsilon^{2}$. If $\mathcal{M}_{\scriptscriptstyle G}$ vanishes for $X\rightarrow 0$ much faster than $X^{1/2}$ the unwanted $\vec{\nabla}\mathcal{M}_{\scriptscriptstyle G}\cdot\vec{\nabla}\epsilon$ term becomes strongly subdominant in the NR limit.
@@ VI Cosmology
Our present picture of cosmology within GR has not emerged as a prediction of the theory – general relativity can, of course, accommodate many other consistent cosmological models. This picture has been built layer upon layer with the benefit of major inputs beyond the adoption of GR: Inputs from physics – as our view of its laws have evolved; inputs from observations - as these have unfolded – e.g., on the realization that galaxies are “island universes”, on the expansion of the universe, etc. The material contents of the universe, with their properties and interactions, are all put in by hand. This includes components – dark matter and dark energy – that we do not even know exist from independent evidence, and whose presence, amounts, distributions, exact nature, and interactions are all free for one to invoke as long as they help square GR with the observations. Symmetries – the cosmological principle – are indicated by observations, but then assumed generally. Initial conditions, and whether, and how they follow from inflation, in any of its many versions, are still moot.
@@
The main constraint on variability of $G$ from cosmology that is relevant to our discussion here comes from big-bang nucleosynthesis (BBN), and emerges essentially from the observed abundances of the light elements, which constrain the expansion rate at the relevant time. Reference [18] discusses such constraints, with the various caveats that underlie them, and concludes: “ … assuming the number of neutrinos to be three, leads to the conclusion that $G$ has not varied from more than 20% since nucleosynthesis. But, allowing for a change both in $G$ and $N\nu$ allows one for a wider range of variation.
@@
In light of the BBN constraint, one would require from an acceptable solution of the theory to yield $\mathcal{M}_{\scriptscriptstyle G}\ll 1$ during BBN, but $\mathcal{M}_{\scriptscriptstyle G}\sim-1$ later on, when “dark-matter” effects are needed to account for the expansion history; in particular, during matter dominance [^11] Without dark matter, matter dominance occurs later than in the standard picture..
The two epochs differ in several regards that could lead to such a variation in the value of $\mathcal{M}_{\scriptscriptstyle G}$: a. The former epoch is radiation dominated, while the latter is essentially matter dominated. b. There may have been differences between the metrics in the two sectors, as the Universe emerges from a period of inflations, e.g., in the form of scalar or tensor (graviton) waves [37, 38]. These could have contributed to the VGMOND  scalar variables, with changes between the two epochs. c. Depending on the initial conditions in the two sectors, the former epoch may have been characterized by very near metric symmetry, hence near vanishing of the scalar variables, while later on asymmetries develop due to independent fluctuations on top of the common FLRW metric, leading to finite values of the scalar variables.
@@
Another possible emergent difference between the two sectors, which I mention only in passing, and which I shall not further pursue here, is in the number of baryons. We do not know how, and when exactly, the amount of baryons in our sector was fixed. The standard assumption is that the universe started with a zero baryon number, and at some point, when the three Sakharov conditions [39] were fulfilled, a very small number imbalance between baryons and antibaryons was created, which is what appears today as the remaining baryons after the vast majority of baryons and antibaryons had annihilated. Depending on the exact mechanism responsible for the slight baryon asymmetry, this small difference could have been different in the two sectors, leading to a difference in the densities of baryons and “twin baryons”.
@@
As a result of the assumed symmetry between the two sectors, they have the same global cosmological evolution, with the space averages of the two metrics remaining the same [^13] Note that the scalar variables do not average to zero, even if the metric difference itself does.. With the standard cosmological principle assumed, the common cosmological metric has the FLWR form.
@@
The subfamily of these, the $X$ scalars of the second type that we employed in the examples of Sec. V.0.4 as arguments of $\mathcal{M}_{\scriptscriptstyle G}$, have $u=0$.
@@
When referring to a “cosmological constant” within the standard model of cosmology, we mean the constant $\Lambda$ that enters the Einstein equation as $G^{\mu\nu}=-8\pi GT^{\mu\nu}-\Lambda g^{\scriptscriptstyle\mu\nu}$, where $T^{\mu\nu}$ is the energy-momentum tensor for matter, and $G$ takes the standard value. If this is not a good description of the real world, possible ambiguity may appear as to what we count as “dark energy” and what as anomalously-behaving matter. The distinction between the contribution of “matter” and cosmological constant is also clear in our VGMOND  case when $\mathcal{M}_{\scriptscriptstyle G}$ and $\tilde{\mathcal{M}}$ are identically constants. Then the field equations for the metrics are
@@
and similarly for $\hat{g}_{\scriptscriptstyle\mu\nu}$. Observing such a behaviour (e.g., in the expansion history), and interpreting it based on GR, we would say that we have $(1+\mathcal{M}_{\scriptscriptstyle G})^{-1}$ times the actual contribution of matter (attributing the excess to dark matter), and identify $\tilde{\mathcal{M}}/2\ell_{\scriptscriptstyle M}^{2}$ as a cosmological constant.
But, generally, when $\mathcal{M}_{\scriptscriptstyle G}$ and $\tilde{\mathcal{M}}$ cannot be treated as constants, the distinction between the extra strength of gravity associated with matter, and the contribution of a “dark-energy” term is not clear cut.
Even within strict BIMOND ($\mathcal{M}_{\scriptscriptstyle G}\equiv 0$), because the metrics appear in the scalar variables, the interaction term, $\tilde{\mathcal{M}}$, gives rise in the field equations to terms that are proportional to $\tilde{\mathcal{M}}g^{\scriptscriptstyle\mu\nu}$ and $\tilde{\mathcal{M}}\hat{g}^{\scriptscriptstyle\mu\nu}$, and so have an effective “equation of state $p=-\rho$, with $\rho$ somewhat space-time dependent. But it also gives rise to terms that are not proportional to $g^{\scriptscriptstyle\mu\nu}$ or $\hat{g}^{\scriptscriptstyle\mu\nu}$ (see Refs. [25, 24]). So, if it is legitimate to take spatial averages of all quantities in the field equation when describing the expansion history we would get dark-energy-like equation of state.

### v2

@@ IV Einstein-Palatini formulation
The Einstein-Palatini formulation of GR starts from the Einstein-Hilbert action $\propto\int g^{1/2}g^{\scriptscriptstyle\mu\nu}R_{\mu\nu}$, where the curvature is derived from a connection ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$, symmetric in $\mu\nu$. In the standard, “metric,” formulation of GR, the connection is taken as the Levi-Civita connection, $\Gamma^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$, derived from the metric – thus making the metric the only gravitational DoF. In the Einstein-Palatini formulation of GR, ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$, as it appears in the expression for the curvature scalar, is taken as an independent gravitational DoF. [^7] A connection appears also in the matter action, via covariant derivatives; there, it is standardly taken as the Levi-Civita, metric connection. Then, the field equations gotten by varying the action over ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$ tell us that for solutions of the theory we do have ${{}^{*}\Gamma}^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}=\Gamma^{\scriptscriptstyle\lambda}_{\scriptscriptstyle\mu\nu}$. Then, variation over the metric gives the standard Einstein equations for the metric, resulting in a theory equivalent to GR.
@@ IV.1 Field equation in the Einstein-Palatini formulation
Expression (31) can be viewed as the (uncoupled) Einstein-Hilbert action for the metrics $\mathfrak{g}_{\scriptscriptstyle\mu\nu}$ and $\hat{\mathfrak{g}}_{\scriptscriptstyle\mu\nu}$. Writing it in this form is the first step in performing the Weyl transformation to the Einstein frame. To complete the transformation we would have to express the metrics and connections appearing in the matter actions and in the BIMOND metric-interaction in terms of $\mathfrak{g}_{\scriptscriptstyle\mu\nu}$, $\hat{\mathfrak{g}}_{\scriptscriptstyle\mu\nu}$, and the starred connections. In principle, such a transformation may lead to a form of the action that is more amenable to scrutiny of the health of the theory. However, it is readily seen that completing the Weyl transformation would be highly involved, and I shall not (and cannot) attempt it here. Some of the complications involved concern the fact that this would lead to matter actions in which matter in each sector would appear to couple to both metrics. We also see, e.g., from Eqs. (32) and (33), that higher derivatives of the metric will enter, via the derivatives of $\mathcal{M}_{\scriptscriptstyle G}$. Whether such a transformation would lead to a more useful form is also questionable; for example, its usefulness may be restricted by possible nonanalyticity features of $\mathcal{M}_{\scriptscriptstyle G}$ (see Sec. V.4). All this is also true for the metric formulation of the theory.
@@ V Observational constraints on $\mathcal{M}_{\scriptscriptstyle G}$ and some examples
Before offering specific examples, I discuss what we require from such theories to achieve in the subcosmological realm, and the resulting constraints on the behavior of $\mathcal{M}_{\scriptscriptstyle G}$.
V.1 Nonrelativistic limit
@@
These facts are used in Sec. V.4 to demonstrate that, with appropriate forms of $\mathcal{M}_{\scriptscriptstyle G}$, the various observational constraints can be easily satisfied.
V.2 High-acceleration limit
@@
This requirement applies to NR systems – such as the inner solar system, and noncompact stars – as well as to relativistic ones – such as black holes, pulsars, and merging compact objects. To achieve it, $\mathcal{M}_{\scriptscriptstyle G}$ has to depend also on variables that do not vanish in the NR limit – variables that are of the first type, but not of the second. We can then make $\mathcal{M}_{\scriptscriptstyle G}$ vanish rapidly as these variables tend to infinity, independently of the values of the variables of the second type on which $\mathcal{M}_{\scriptscriptstyle G}$ depends.
V.3 Low-acceleration limit
@@
V.4 Example
To summarize, I want to demonstrate that $\mathcal{M}_{\scriptscriptstyle G}$ can be chosen so that the following requirements are satisfied: (1) $\mathcal{M}_{\scriptscriptstyle G}$ vanishes rapidly for all high-acceleration systems, relativistic and NR, to comply with the many constraints from subcosmological systems on anomalous $G$ values. (2) $\mathcal{M}_{\scriptscriptstyle G}$ also vanishes rapidly in the NR limit even for systems that are not high-acceleration, so that pure-BIMOND phenomenology is predicted for galactic systems. (3) $\mathcal{M}_{\scriptscriptstyle G}$ can take finite values in later-time cosmology, while it vanishes for early-time cosmology. In this section I discuss the fulfilment of requirements (1) and (2) for subcosmological systems. Cosmology will be discussed in Sec. VI.
@@ V.4 Example
As we saw in Sec. V.1, for relativistic, subcosmological systems, of mass $M$, and characteristic size $\ell$, both the $X$ and $Y$ variables are of order $(MG/\ell^{2}a_{0})^{2}\ggg 1$ (because they are perforce of high accelerations). [^10] For highly relativistic systems, such as the vicinity or a black-hole horizon, $X,Y\sim\ell_{\scriptscriptstyle U}/\ell$, where $\ell_{\scriptscriptstyle U}$ is the Hubble distance. For NR, subcosmological systems, $Y=O(g^{2}/a_{0}^{2})$, where $g=MG/\ell^{2}$ for high-acceleration systems, and $g=(MGa_{0}/\ell^{2})^{1/2}$ for low-acceleration ones. And, $X=O(Y\epsilon^{2})$. In the low-acceleration case ($g\not\gg a_{0}$), we have $X\lesssim\epsilon^{2}\lll 1$. To ensure that requirement (2) is satisfied we need $\mathcal{M}_{\scriptscriptstyle G}$ to vanish rapidly enough in the low-acceleration case, as $X\lesssim\epsilon^{2}\rightarrow 0$.
To see how fast, consider the NR limit of the Einstein-Hilbert Lagrangian. We saw in Eqs. (29)-(30) that the post-Newtonically dominant term is $\vec{\nabla}\mathcal{M}_{\scriptscriptstyle G}\cdot\vec{\nabla}\epsilon$, which we want to be negligible compared with the Poissonian term $(\vec{\nabla}\epsilon)^{2}$. Since $X\lesssim\epsilon^{2}$. If $\mathcal{M}_{\scriptscriptstyle G}$ vanishes for $X\rightarrow 0$ much faster than $X^{1/2}$ the unwanted $\vec{\nabla}\mathcal{M}_{\scriptscriptstyle G}\cdot\vec{\nabla}\epsilon$ term becomes strongly subdominant in the NR limit.
@@
The expression for $\mathcal{M}_{\scriptscriptstyle G}$ in the example above is nonanalytic in $Y$ in the high-acceleration limit – it cannot be expanded in powers of $a_{0}$ near the GR limit $a_{0}=0$. It is also nonanalytic in the variable $X$ near the decoupling limit $X=0$ (also a GR limit with a different CC). This nonanalyticity can stand in the way of applying some standard techniques to assess the health of the theory. But this, in itself, does not argue against such a behavior. In quantum theory, for example, we encounter expression – such as the barrier-penetration probability, or the path weight in the path-integral approach – that have similar nonanalytic dependence on $\hbar$ at the classical limit $\hbar=0$. (MOND is also nonanalytic in the deep-MOND limit due to the appearance of fractional powers in the Lagrangian – forced by phenomenology.)
In fact, I choose this example behavior intentionally for two reasons. The first is that phenomenological constraints – in particular solar-system constraints – require that the GR limit is approached very fast in the high-acceleration limit ($a_{0}\rightarrow 0$), in $\mathcal{M}_{\scriptscriptstyle G}$ as in $\tilde{\mathcal{M}}$. Also, I wanted an example of $\mathcal{M}_{\scriptscriptstyle G}$ that vanishes fast for small $X$, so as to be compatible with BBN limits, on one hand, and, on the other, to grow rapidly to its asymptotic constant value, so as to account for faux-DM effects in the expansion history. All this can be achieved perhaps with analytic functions, which, however, I do not see as a pressing desideratum. The second, and stronger, reason is that there exist general theorems to the effect that bimetric theories, generically, possess unwanted ghost instabilities (see, e.g., Refs. [37, 38, 30]). However, such theorems build on expansions in the metric coupling near the decoupling limit; so, opting for nonanalytic behavior may escape the sentence of such theorems, even though, in itself, it does not guarantee the absence of ghosts. (See also the comments on the possible presence of ghosts in VGMOND and in BIMOND generally, in Sec. I.)
@@ VI Cosmology
Our present picture of cosmology within GR has not emerged as a prediction of the theory – general relativity can, of course, accommodate many other consistent cosmological models. This picture has been built layer upon layer with the benefit of major inputs beyond the adoption of GR: Inputs from physics – as our view of its laws have evolved; inputs from observations – as these have unfolded – e.g., on the realization that galaxies are “island universes,” on the expansion of the universe, etc. The material contents of the universe, with their properties and interactions, are all put in by hand. This includes components – dark matter and dark energy – that we do not even know exist from independent evidence, and whose presence, amounts, distributions, exact nature, and interactions are all free for one to invoke as long as they help square GR with the observations. Symmetries – the cosmological principle – are indicated by observations, but then assumed generally. Initial conditions, and whether, and how they follow from inflation, in any of its many versions, are still moot.
@@
The main constraint on variability of $G$ from cosmology that is relevant to our discussion here comes from big bang nucleosynthesis (BBN), and emerges essentially from the observed abundances of the light elements, which constrain the expansion rate at the relevant time. Reference [18] discusses such constraints, with the various caveats that underlie them, and concludes: “ … assuming the number of neutrinos to be three, leads to the conclusion that $G$ has not varied from more than 20% since nucleosynthesis. But, allowing for a change both in $G$ and $N\nu$ allows one for a wider range of variation.
@@
In light of the BBN constraint, one would require from an acceptable solution of the theory to yield $\mathcal{M}_{\scriptscriptstyle G}\ll 1$ during BBN, but $\mathcal{M}_{\scriptscriptstyle G}\sim-1$ later on, when “dark matter” effects are needed to account for the expansion history; in particular, during matter dominance [^11] Without dark matter, matter dominance occurs later than in the standard picture..
The two epochs differ in several regards that could lead to such a variation in the value of $\mathcal{M}_{\scriptscriptstyle G}$: (a) The former epoch is radiation dominated, while the latter is essentially matter dominated. (b) There may have been differences between the metrics in the two sectors, as the Universe emerges from a period of inflations, e.g., in the form of scalar or tensor (graviton) waves [40, 41]. These could have contributed to the VGMOND  scalar variables, with changes between the two epochs. (c) Depending on the initial conditions in the two sectors, the former epoch may have been characterized by very near metric symmetry, hence near vanishing of the scalar variables, while later on asymmetries develop due to independent fluctuations on top of the common FLRW metric, leading to finite values of the scalar variables.
@@
Another possible emergent difference between the two sectors, which I mention only in passing, and which I shall not further pursue here, is in the number of baryons. We do not know how, and when exactly, the amount of baryons in our sector was fixed. The standard assumption is that the universe started with a zero baryon number, and at some point, when the three Sakharov conditions [42] were fulfilled, a very small number imbalance between baryons and antibaryons was created, which is what appears today as the remaining baryons after the vast majority of baryons and antibaryons had annihilated. Depending on the exact mechanism responsible for the slight baryon asymmetry, this small difference could have been different in the two sectors, leading to a difference in the densities of baryons and “twin baryons.”
@@
As a result of the assumed symmetry between the two sectors, they have the same global cosmological evolution, with the space averages of the two metrics remaining the same [^13] Note that the scalar variables do not average to zero, even if the metric difference itself does.. With the standard cosmological principle assumed, the common cosmological metric has the FLRW form.
@@
The subfamily of these, the $X$ scalars of the second type that we employed in the examples of Sec. V.4 as arguments of $\mathcal{M}_{\scriptscriptstyle G}$, have $u=0$.
@@
When referring to a “cosmological constant” within the standard model of cosmology, we mean the constant $\Lambda$ that enters the Einstein equation as $G^{\mu\nu}=-8\pi GT^{\mu\nu}-\Lambda g^{\scriptscriptstyle\mu\nu}$, where $T^{\mu\nu}$ is the energy-momentum tensor for matter, and $G$ takes the standard value. If this is not a good description of the real world, possible ambiguity may appear as to what we count as dark energy and what as anomalously-behaving matter. The distinction between the contribution of matter and cosmological constant is also clear in our VGMOND  case when $\mathcal{M}_{\scriptscriptstyle G}$ and $\tilde{\mathcal{M}}$ are identically constants. Then the field equations for the metrics are
@@
and similarly for $\hat{g}_{\scriptscriptstyle\mu\nu}$. Observing such a behavior (e.g., in the expansion history), and interpreting it based on GR, we would say that we have $(1+\mathcal{M}_{\scriptscriptstyle G})^{-1}$ times the actual contribution of matter (attributing the excess to dark matter), and identify $\tilde{\mathcal{M}}/2\ell_{\scriptscriptstyle M}^{2}$ as a cosmological constant.
But, generally, when $\mathcal{M}_{\scriptscriptstyle G}$ and $\tilde{\mathcal{M}}$ cannot be treated as constants, the distinction between the extra strength of gravity associated with matter, and the contribution of a dark energy term is not clear cut.
Even within strict BIMOND ($\mathcal{M}_{\scriptscriptstyle G}\equiv 0$), because the metrics appear in the scalar variables, the interaction term, $\tilde{\mathcal{M}}$, gives rise in the field equations to terms that are proportional to $\tilde{\mathcal{M}}g^{\scriptscriptstyle\mu\nu}$ and $\tilde{\mathcal{M}}\hat{g}^{\scriptscriptstyle\mu\nu}$, and so have an effective “equation of state $p=-\rho$, with $\rho$ somewhat space-time dependent. But it also gives rise to terms that are not proportional to $g^{\scriptscriptstyle\mu\nu}$ or $\hat{g}^{\scriptscriptstyle\mu\nu}$ (see Refs. [25, 24]). So, if it is legitimate to take spatial averages of all quantities in the field equation when describing the expansion history we would get an equation of state that describes dark energy.